Using this function W(d)=0.112d approximates the amount of water in cm that results from d of snow melting, what would be the depths of 15cm,21cm,72cm. Can you help? Thank you! Shelly
Answer by stanbon(75887)

Using this function W(d)=0.112d approximates the amount of water in cm that results from d of snow melting, what would be the depths of 15cm,21cm,72cm. Can you help?
------
W(15cm) = 0.112(15) = 1.68
W(21cm) = 0.112(21) = 2.352
W(72cm) = 0.112(72) = 8.064
===============================
Cheers,
Stan H.